### v2

# Exploring Prompting Methods for Mitigating Class Imbalance through Synthetic Data Generation with Large Language Models

Abstract
Large language models (LLMs) have demonstrated impressive in-context learning capabilities across various domains. Inspired by this, our study explores the effectiveness of LLMs in generating realistic tabular data to mitigate class imbalance. We investigate and identify key prompt design elements such as data format, class presentation, and variable mapping to optimize the generation performance. Our findings indicate that using CSV format, balancing classes, and employing unique variable mapping produces realistic and reliable data, significantly enhancing machine learning performance for minor classes in imbalanced datasets. Additionally, these approaches improve the stability and efficiency of LLM data generation. We validate our approach using six real-world datasets and a toy dataset, achieving state-of-the-art performance in classification tasks. The code is available here.
Figure 1: Generation results on an imbalanced toy dataset. The major class is shown in blue, and the minor class in red. Our approach via in-context learning of LLM achieves (1) clear class boundaries, (2) accurate feature correlations, (3) proper value ranges, (4) well-represented relationships between categorical and numerical variables (last row), and (5) better diversity compared to the fine-tuned GReaT [4] model. Detailed explanations and complete results are provided in Appendix E.
1 Introduction
Tabular data plays a crucial role in various classification tasks, such as disease diagnosis [1] and financial analysis [3, 24]. However, challenges such as data scarcity and class imbalance, particularly for rare yet crucial events, can severely degrade the performance of machine learning (ML) models. To address these, synthetic data generation methods have been developed to augment tabular datasets cost-effectively. Traditional methods such as SMOTE and its variants [6, 12, 20] focus on generating data for minority classes to alleviate class imbalance. Recently, advanced deep learning-based generation models have shown promising results, such as TVAE [30], CTAB-GAN [33], and TabDDPM [18].
Large language models (LLMs) have also demonstrated significant potential in generating realistic tabular data, showing robust performance across numerical and categorical data [4, 31]. However, these models often require extensive, data-specific fine-tuning, which can lead to overfitting on major classes or dominant feature values when applied to small and imbalanced tabular datasets. An alternative approach leverages the in-context learning capabilities of LLMs. Extensive studies have demonstrated that LLMs can solve complex problems, learn and mimic patterns from input data, and generate new data without parameter updates [13, 17, 19, 28, 29, 32]. This capability holds significant potential for generating realistic and reliable tabular data to address class imbalance.
However, the in-context learning capabilities of LLMs largely depend on a prompt design [10, 26, 28, 32], while these methods are not directly applicable to tabular data generation, especially for imbalanced class scenarios. This is because tabular data is not inherently in plain text and involves unique challenges such as identifying feature correlations and accurately representing underrepresented attributes. Therefore, carefully crafted prompt methods specifically for imbalanced tabular data are essential to maximize LLM capabilities in generating realistic data effectively. While several studies have highlighted the potential of LLMs for generating synthetic tabular data using in-context learning [25, 31], they have not sufficiently explored various prompt design elements, such as data format or class presentation. This study aims to investigate the effectiveness of LLMs in resolving class imbalance problems and identify the optimal prompts to maximize this effectiveness.
Figure 2: Comparison of ML classification performance when adding synthetic data to the Travel dataset. The results are averaged across four classifiers, XGBoost, CatBoost, LightGBM, and Gradient Boosting Classifier. Each classifier is run five times, resulting in a total of 20 experiments.
To answer these questions, we explore various prompt methods and analyze the generated results to identify specific prompt structures and templates that significantly improve the quality of synthetic tabular data generated by LLMs, thereby enhancing ML classification performance. We propose several prompt design elements, including data format, class presentation methods, variable mapping, and task specification. Our findings indicate that providing data in CSV format, balancing and grouping classes, and applying a unique variable mapper to distinguish different variables significantly improve the quality and efficiency of synthetic data generation, underscoring their effectiveness in optimizing LLM performance. As shown in Fig. 1, the data generated using this prompt design exhibit the most distinct boundaries between target classes, accurate feature correlation, and variable ranges.
Moreover, this approach significantly improves the classification performance of ML models when applied to real-world datasets. As shown in Fig. 2, baseline methods exhibit low sensitivity, indicating that they fail to accurately generate minority class samples due to inherent class biases. In contrast, our method overcomes these, achieving high sensitivity and balanced performance for all classes.
In summary, our contributions are:
  • Our study explores and analyzes the effectiveness of LLMs in generating realistic synthetic tabular data via in-context learning, offering guidelines for designing prompts tailored to effectively exploit LLMs for this task, particularly addressing class imbalance.
  • We propose simple yet effective prompting methods that generate high-quality class-specific tabular data, maintaining feature correlation and value ranges.
  • The proposed approach significantly improves ML classification performance and enhances data generation efficiency and stability via in-context learning.
  • The proposed approach is generally applicable and easy to implement for any tabular classification task with minimal pre-processing, providing publicly available code.
  • We validate the effectiveness of our approach through extensive evaluation across six real-world public tabular data classification datasets and a toy dataset.
Figure 3: Overview of our approach. Our prompt includes repeated data example sets consisting of feature names and class-balanced groups, with the feature name at the end serving as a trigger for the LLM to generate realistic synthetic tabular data. The proposed unique variable remaps categorical values to distinct alphanumeric strings, ensuring clear distinction and variability among variables.
2 Method
In this study, we explore various prompt design options to maximize the in-context learning capability of LLMs for generating high-quality tabular data. Our goal is to propose an optimal approach that reliably and efficiently produces realistic tabular data to enhance ML classification performance, especially in the context of class imbalance. We start with a tabular dataset $T\in\mathds{R}^{n\times m}$, comprising $n$ instances, each with $m$ variables. We aim to generate a synthetic dataset, $\hat{T}\in\mathds{R}^{n^{\prime}\times m}$, which accurately mimics the characteristics of $T$, including class conditions, feature correlations, and value ranges. To achieve this, we explore several key prompt design elements: Data format, class presentation methods, variable mapping, and task specification.
We evaluate these design choices based on the quality of the generated data, with a primary focus on ML classification performance and generation efficiency, as detailed in Section 3. Our carefully crafted prompt, as depicted in Fig.3, begins with optional variable descriptions followed by multiple sets of data samples. Each set consists of feature names and samples grouped by target class. By adding feature names at the end of the prompt as a trigger, the LLM generates realistic data that improves ML performance and maintains feature correlations similar to the original data. Prompt examples and an illustration of the ablated versions are provided in Appendices I.5 and I.1, respectively.
2.1 Data format: Sentence vs CSV style
We investigate effective data formatting methods to represent tabular data within prompts, ensuring that LLMs can interpret the demonstrations and generate new data accurately. Tabular data can be formatted as plain text or as comma-separated values (CSV style). Let the feature name of the $k$-th variable be $v_{k}$, and the $i$-th observation of this variable be $o_{k,i}$. According to previous work [4], this value can be transformed into a sentence-style representation as $[v_{k},``is",o_{k,i},``,"]$. In contrast, the CSV style can be expressed as $[o_{k,i},``,"]$, presenting only values, with variable names for all variables specified as $\bm{v}=[v_{1},``,",v_{2},``,",\cdots,v_{m}]$ at the beginning of the data sample presentation. Here, the sentence style redundantly uses variable names and “is" tokens for every value, leading to higher token usage and computational cost than the CSV style. The CSV-style format allows for a higher volume of in-context learning examples within the same token constraints. Thus, our method utilizes a CSV-style format to maximize the number of examples, as providing a large number of data samples is crucial for ensuring sufficient representation of the original dataset in in-context learning.
2.2 Class presentation
Class presentation methods aim to ensure that all target classes are adequately represented within a prompt, particularly addressing class imbalance for minority classes.
Single-class vs. Multi-class generation    When generating data with class conditions, a primary consideration is whether to generate data for one class at a time or for all classes simultaneously. Generating data for a single class allows the model to focus on specific class characteristics, whereas generating data for multiple classes simultaneously enables comparison and contrast across classes. Our experiments indicate that generating data for multiple classes simultaneously, in a structured manner (explained subsequently), produces samples that more accurately represent each class compared to single-class generation. This suggests that generating data with an awareness of the characteristics of other classes is more effective.
Original vs. Balanced class ratio    When constructing a prompt to generate multiple classes simultaneously, naive random sampling from the original tabular dataset $T$ often results in an imbalanced ratio of samples for each target class. To address this, we consider balancing the sample numbers for minor and major classes rather than adhering to the original data distribution. Our findings indicate that balancing the sample numbers significantly improves the quality of generated data for minor classes, demonstrating that training with balanced data helps LLMs learn the true distributions of target classes more effectively.
Listing vs. Class grouping    Within a prompt, the data examples can either be listed in the order they are sampled without class distinction or grouped by class. Empirical evidence suggests that grouping by class leads to a better representation of class-specific feature correlations in the generated data. This grouping emphasizes contrasts between different classes and similarities within each group. As a result, the LLM generates more distinct samples for each class and more coherent samples within each class. Additionally, group-wise prompting can effectively condition LLMs to generate data with specific values. Without grouping, it is challenging to constrain LLMs to generate samples with specific attributes, often requiring extensive iterations until the desired samples are produced, which is inefficient. However, when examples are provided in groups by class, LLMs learn this and generate samples for each class in groups. Thus, grouping is crucial for conditioning the LLM to generate data with specific values and conditions. Moreover, grouping can be extended to condition on any arbitrary features other than target classes by grouping samples based on specific requirements.
2.3 Variable mapping: Original values vs. Unique variable mapping
As with other LLM-based approaches [4], our approach allows minimal data preprocessing, maintaining the original feature names and values. This significantly reduces the need for complex data manipulations often required in existing methods [18, 34] and helps to preserve the integrity of the raw data. However, one critical consideration arises when a dataset includes a large number of categorical variables that share the same values (e.g., extensive boolean variables). In this case, the data examples in the prompt can become monotonous and repetitive, making it challenging for the model to distinguish between different variables. This can significantly degrade the quality of the generated data. Moreover, when the LLM receives such input, the number of data samples generated at once can significantly decrease, and the generation efficiency can greatly decline.
To address this, we propose a unique variable mapping method, which remaps the values of each categorical variable to distinct random alphanumeric strings, as depicted in Fig. 3. Uniform substitution is applied across the dataset, maintaining the integrity of the data structure while introducing necessary variation. For example, consider a categorical variable $v_{1}$ with values of ‘t’ and ‘f’. Our approach remaps these values to distinct three-character strings, such as ‘JY0’ and ‘GGN,’ respectively, and applies them consistently across all data points in the dataset. Other categorical variables are similarly remapped to unique values. The transformed dataset is then used as in-context learning demonstrations. Although these new values do not hold inherent semantic meaning, they effectively represent categorical distinctions as symbols, allowing the LLM to identify and utilize patterns within them. This approach ensures that each variable has a unique representation, making the variables clearly distinguishable and contributing to more accurate and efficient data generation.
2.4 Task specification
A crucial aspect of the prompt design is to clearly define the task within the prompt, guiding the LLM to generate synthetic data as intended. To achieve this, the following options can be considered:
Explicit instruction vs. Completion triggering    We can clarify the task to the LLM by including explicit instructions to “generate new data samples." However, crafting specific instructions can be challenging, as LLMs are sensitive to minor variations in the prompt, making this approach complex for the tabular data generation task. A simpler and more effective method is providing input patterns that LLM can mimic to generate new data, only using data samples without explicit instruction. For example, let the $i$-th set of data samples for all $c$ classes be formatted as $\bm{S}_{i}=[\bm{v},\bm{s}_{i,1},\cdots,\bm{s}_{i,c}]$, where $\bm{s}_{i,j}$ consists of $n$ data samples for the $j$-th class. Here, the feature name $\bm{v}$ works as a header to indicate the start of a set. We input $t$ such sets in a prompt, $[\bm{S}_{1},\bm{S}_{2},\ldots,\bm{S}_{t}]$, creating a pattern and expecting the LLM to generate $\bm{S}_{t\texttt{+}1}=[\bm{v},\bm{s}_{t\texttt{+}1,1},\cdots,\bm{s}_{t\texttt{+}1,c}]$. To ensure this, we add a trigger to induce the LLM to complete the sequence. That is, we include the header of the new set $[\bm{v}]$ at the end of the prompt, which cues the LLM to generate the rest, $[\bm{s}_{t\texttt{+}1,1},\cdots,\bm{s}_{t\texttt{+}1,c}]$, learning the pattern from the input data that these groups are followed by $\bm{v}$. Thus, by creating a pattern in the input prompt, we can condition the LLM to generate a specific number ($c\times n$) of samples for each target class.
Providing contextual information    Optionally, including contextual information in the prompt can help the LLM understand the data structure and generate accurate samples. Therefore, if available, we add line-by-line variable descriptions at the beginning of the prompt. Additionally, one can consider clarifying the target class values at the beginning of each group. However, we observe that treating the target variable as one of the attributes and placing it as the first variable yields better performance, helping the LLM generate more coherent and accurate samples.
2.5 Overview of tabular data generation process
Synthetic data generation starts with creating a prompt template (see Appendix I.5 for examples). This template determines the target class and the number of sets, groups, and samples per group to construct demonstrations. If numerous categorical variables share the same unique values, we apply unique variable mapping to the dataset. Next, we randomly sample data examples from the dataset to fill in this template. Based on the provided examples, the LLM generates new data, which is then converted back into a tabular format. Instances with categorical variable values not present in the training data are discarded to maintain data authenticity. This process is repeated until the desired number of samples is obtained. Each iteration involves newly sampled examples from the original dataset, allowing the LLM to encounter various combinations and produce diverse samples. Although each individual generation is based on the distribution of the given subset, cumulatively, the iterations enable the model to explore the entire range of the original data. This method ensures that the output generated by the LLM comprehensively reflects the original data characteristics (See Appendix G).
Table 1: Comparison of machine learning classification performance when synthetic data are added to the original data. The results are averaged across four classifiers, with each model run five times. Complete results, including all baselines and standard deviation values, are provided in Appendix K.
Dataset Method #syn F1 score $\uparrow$ BAL ACC $\uparrow$ Sensitivity $\uparrow$ Specificity $\uparrow$
Travel Original - 58.12 (0.00) 71.00 (0.00) 57.00 (0.00) 85.00 (0.00)
+TVAE [30] +1K 59.78 (+1.66) 72.35 (+1.35) 62.00 (+5.00) 82.69 (-2.31)
+CTAB-GAN+ [34] +1K 54.66 (-3.46) 68.62 (-2.38) 53.00 (-4.00) 84.23 (-0.77)
+GReaT [4] +1K 60.95 (+2.83) 72.86 (+1.86) 58.80 (+1.80) 86.92 (+1.92)
+TabDDPM [18] +1K 53.20 (-4.92) 67.70 (-3.30) 50.40 (-6.60) 85.00 (0.00)
+Ours +1K  66.65 (+8.53) 78.23 (+7.23) 78.00 (+21.00) 78.46 (-6.54)
Sick Original - 87.81 (0.00) 91.22 (0.00) 82.83 (0.00) 99.61 (0.00)
+TVAE [30] +1K 87.77 (-0.04) 91.47 (+0.25) 83.37 (+0.54) 99.56 (-0.05)
+CTAB-GAN+ [34] +1K 82.35 (-5.46) 86.28 (-4.94) 72.83 (-10.00) 99.74 (+0.13)
+GReaT [4] +1K 87.23 (-0.58) 90.83 (-0.39) 82.07 (-0.76) 99.60 (-0.01)
+TabDDPM [18] +1K 85.17 (-2.64) 89.30 (-1.92) 79.02 (-3.81) 99.57 (-0.04)
+Ours +1K  88.71 (+0.90) 92.93 (+1.71) 86.41 (+3.58) 99.44 (-0.17)
HELOC Original - 71.01 (0.00) 73.21 (0.00) 67.89 (0.00) 78.52 (0.00)
+TVAE [30] +1K 71.12 (+0.11) 73.25 (+0.04) 68.15 (+0.26) 78.34 (-0.18)
+CTAB-GAN+ [34] +1K 71.03 (+0.02) 73.15 (-0.06) 68.13 (+0.24) 78.17 (-0.35)
+GReaT [4] +1K 70.35 (-0.66) 72.96 (-0.25) 66.22 (-1.67) 79.70 (+1.18)
+TabDDPM [18] +1K 70.65 (-0.36) 72.89 (-0.32) 67.51 (-0.38) 78.26 (-0.26)
+Ours +1K  71.92 (+0.91) 73.66 (+0.45) 69.96 (+2.07) 77.35 (-1.17)
Income Original - 66.90 (0.00) 76.45 (0.00) 57.28 (0.00) 95.61 (0.00)
+TVAE [30] +20K 66.96 (+0.06) 76.80 (+0.35) 59.13 (+1.85) 94.48 (-1.13)
+CTAB-GAN+ [34] +20K 66.49 (-0.41) 76.42 (-0.03) 58.14 (+0.86) 94.70 (-0.91)
+GReaT [4] +20K 67.95 (+1.05) 77.51 (+1.06) 60.69 (+3.41) 94.33 (-1.28)
+TabDDPM [18] +20K 66.85 (-0.05) 76.50 (+0.05) 57.70 (+0.42) 95.30 (-0.31)
+Ours +20K  69.16 (+2.26) 79.15 (+2.70) 66.45 (+9.17) 91.85 (-3.76)
Diabetes Original - 54.87 (0.00) 42.07 (0.00) 60.00 (0.00) 60.73 (0.00)
+TVAE [30] +10K 54.79 (-0.08) 41.96 (-0.11) 59.96 (-0.04) 60.71 (-0.02)
+CTAB-GAN+ [34] +10K 54.24 (-0.63) 41.52 (-0.55) 59.63 (-0.37) 60.01 (-0.72)
+GReaT [4] +10K 54.78 (-0.09) 41.98 (-0.09) 59.98 (-0.02) 60.61 (-0.12)
+TabDDPM [18] +10K 54.64 (-0.23) 41.83 (-0.24) 59.91 (-0.09) 60.55 (-0.18)
+Ours +10K  54.94 (+0.07) 42.14 (+0.07) 60.04 (+0.04) 60.82 (+0.09)
Thyroid Original - 94.23 (0.00) 95.08 (0.00) 91.14 (0.00) 99.02 (0.00)
+TVAE [30] +1K 90.45 (-3.78) 92.20 (-2.88) 86.36 (-4.78) 98.04 (-0.98)
+CTAB-GAN+ [34] +1K 27.46 (-66.77) 58.07 (-37.01) 16.14 (-75.00) 100.0 (+0.98)
+GReaT [4] +1K 91.31 (-2.92) 92.46 (-2.62) 85.91 (-5.23) 99.02 (0.00)
+TabDDPM [18] +1K 94.39 (+0.16) 96.26 (+1.18) 95.45 (+4.31) 97.06 (-1.96)
+Ours +1K  94.80 (+0.57) 96.39 (+1.31) 95.23 (+4.09) 97.55 (-1.47)
3 Experiments
Datasets    We use six real-world classification datasets from diverse domains to evaluate our method: Travel, Sick, HELOC, Income, Diabetes, and Thyroid. The Thyroid dataset is released after the training cut-off date for GPT-3.5-turbo-0613, which provides a more rigorous validation of our approach on the completely unseen dataset. For binary classification datasets, the minor class is set to be one. All duplicate data samples are removed from the datasets. Subsequently, we randomly split them into 80% training and 20% test sets. More details for the datasets are provided in Appendix I.2.
Evaluation measure    We evaluate our generation methodology from three critical aspects: ML classification performance, data generation success rate and stability with LLM in-context learning, and validity of generated data in terms of feature correlation and value range. For the classification task, we measure the F1 score, sensitivity, specificity, and balanced accuracy (BAL ACC). To measure the feature correlation, we use Pearson correlation and Cramér’s V correlation.
Baseline models    In our study, we compare our method with various generative models for tabular data generation, including TVAE [30], CopulaGAN [21], CTGAN [30], CTAB-GAN [33], CTAB-GAN+ [34], GReaT [4], and TabDDPM [18]. This comprehensive evaluation assesses the effectiveness of our approach across different data synthesis paradigms. To compare the prompting method, we also use the prompts used in CuratedLLM [25] and LITO [31].
Experimental details    Our method utilizes the OpenAI GPT-3.5 models, the GPT-3.5-turbo-0613 and GPT-3.5-turbo-16k-0613 versions, Mistral-7b-v0.1 [14], and Llama-2-7b [27]. Unless stated otherwise, four machine learning classification models are used: XGBoost [8], CatBoost [23], LightGBM [16], and Gradient boosting classifier [11]. We execute five independent runs for each model and report the average values. Additional details are provided in Appendix I.
Table 2: Comparison of classification performance using synthetic data generated by open-source LLMs, Llama2 and Mistral, with the proposed prompting strategy, utilizing the Gradient Boosting classifier. We report the average performances for five runs. #syn denotes the number of synthetic samples integrated into the original dataset.
Dataset Method Model #syn F1 score $\uparrow$ BAL ACC $\uparrow$ Sensitivity $\uparrow$ Specificity $\uparrow$
Travel Original - - 60.00$\pm$0.00 72.31$\pm$0.00 60.00$\pm$0.00 84.62$\pm$0.00
+Ours Mistral +1K 66.67$\pm$0.00 78.00$\pm$0.00 76.00$\pm$0.00 80.00$\pm$0.00
+Ours Llama2 +1K 67.80$\pm$0.00 79.23$\pm$0.00 80.00$\pm$0.00 78.46$\pm$0.00
+Ours GPT3.5 +1K 70.18$\pm$0.00 80.77$\pm$0.00 80.00$\pm$0.00 81.54$\pm$0.00
Sick Original - - 84.09$\pm$0.00 89.86$\pm$0.00 80.43$\pm$0.00 99.28$\pm$0.00
+Ours Mistral +1K 88.42$\pm$0.00 95.15$\pm$0.00 91.30$\pm$0.00 99.00$\pm$0.00
+Ours Llama2 +1K 85.53$\pm$0.95 93.14$\pm$0.52 87.39$\pm$0.97 98.88$\pm$0.06
+Ours GPT3.5 +1K 91.95$\pm$0.00 93.41$\pm$0.00 86.96$\pm$0.00 99.86$\pm$0.00
HELOC Original - - 71.32$\pm$0.04 73.39$\pm$0.03 68.47$\pm$0.06 78.31$\pm$0.00
+Ours Mistral +1K 71.48$\pm$0.00 73.74$\pm$0.00 68.00$\pm$0.00 79.47$\pm$0.00
+Ours Llama2 +1K 70.77$\pm$0.00 73.08$\pm$0.00 67.37$\pm$0.00 78.79$\pm$0.00
+Ours GPT3.5 +1K 71.93$\pm$0.02 73.68$\pm$0.02 69.90$\pm$0.00 77.45$\pm$0.04
3.1 Machine learning classification performance using the synthetic data
We assess the quality of the synthetic data samples by comparing the machine learning classification performance averaged over 20 experiments. As shown in Table 1, our method achieves state-of-the-art F1 scores and balanced accuracy across all six datasets, surpassing the baselines that require model training. Notably, our method is the only one that consistently outperforms the original data in both F1 score and balanced accuracy across all six datasets. Other methods often result in worse performance than the original data, indicating that the data generated by baseline methods disrupts the original data distribution. These findings underscore the robustness and effectiveness of our method, affirming its superiority over baseline models, even for challenging datasets.
Moreover, baseline methods such as GReaT and TabDDPM exhibit significantly lower sensitivity compared to specificity, particularly for Travel and Income. Their high balanced accuracy is due to high specificity, but their severely low sensitivity indicates a failure to generate appropriate minority class data. In contrast, our method significantly improves sensitivity and achieves a balance between sensitivity and specificity while also attaining the best-balanced accuracy and F1 score compared to the baselines. For example, in Travel, our method achieves a sensitivity of 78%, which is 19%p higher than the second-best method, GReaT, at 58.8%. In Income, our method shows 66.45% sensitivity, surpassing the second-best score of 60.69%. These results highlight the effectiveness of our method in accurately representing minority classes in tabular datasets.
Overall, our findings demonstrate that our approach is generally applicable across various imbalanced tabular datasets and excels in generating high-quality samples that improve ML classification performance. For detailed confusion matrix results, see Appendix J, and for further analyses exploring the impact of adding synthetic data only to the minor class or using synthetic data alone, see Appendix A.
3.2 Open-source LLMs
We also apply our prompting method to open-source LLMs, including Llama2 [27] and Mistral [14]. As shown in Table 2, our method exhibits robust performance when used with open-source LLMs across various datasets. Notably, in the Sick dataset, Mistral demonstrates the highest balanced accuracy, indicating its effectiveness in generating high-quality tabular data. These results validate the broader applicability and general effectiveness of our approach with open-source LLMs.
Table 3: Comparison of F1 scores with ablated methods using the Gradient Boosting classifier. Asterisk (*) in Income indicates where only $\texttt{+}1$K synthetic samples are used due to a low success rate in sample generation. The ‘-’ symbol denotes where our unique variable mapping is not needed.
Format Class Balance Group Unique Sick ($\texttt{+}1$K) Travel ($\texttt{+}1$K) Income ($\texttt{+}20$K)
Sentence single 81.52$\pm$0.65 43.24$\pm$0.00 69.62$\pm$0.00
CSV-style single 86.60$\pm$0.00 58.67$\pm$0.00 70.49$\pm$0.00
Sentence multi 77.95$\pm$0.33 58.33$\pm$0.00 67.96$\pm$0.00
CSV-style multi 87.74$\pm$0.96 50.98$\pm$0.00 68.56$\pm$0.00
Sentence multi 81.68$\pm$0.49 55.56$\pm$0.00 69.26$\pm$0.00
CSV-style multi 84.68$\pm$0.38 70.37$\pm$0.00 70.64$\pm$0.00
CSV-style multi 86.60$\pm$0.00 - -
CSV-style multi 81.55$\pm$0.00 70.18$\pm$0.00 70.17$\pm$0.00
CSV-style multi 91.95$\pm$0.00 - -
Table 4: Comparison of token efficiency and generation stability on the Income dataset.
Format Class Balance Group Unique #set Input tokens $\downarrow$ Output samples $\uparrow$ Success rate $\uparrow$
Sentence single - 1832.9 1.26 96%
CSV-style single - 1024.6 1.11 16%
Sentence multi - 1827.4 0.97 97%
CSV-style multi - 912.8 0.38 38%
Sentence multi - 1938.7 1.73 97%
CSV-style multi - 920.3 0.88 31%
CSV-style multi 1 965.6 8.6 48%
CSV-style multi 2 1014.6 10.52 94%
Table 5: Comparison of classification performance on task specification elements in prompt design, averaged across four classifiers: XGBoost, CatBoost, LightGBM, and Gradient Boosting classifier. Methods marked with an asterisk (*) use prompts as proposed in the respective papers.
Method #syn F1 score $\uparrow$ BAL ACC $\uparrow$ Sensitivity $\uparrow$ Specificity $\uparrow$
Instruction-$\text{CuratedLLM}^{*}$ +1K 15.86$\pm$5.13 56.02$\pm$3.54 17.17$\pm$11.80 94.86$\pm$4.93
Instruction-$\text{LITO}^{*}$ +1K 21.32$\pm$1.65 72.06$\pm$2.49 84.46$\pm$3.10 59.66$\pm$3.50
Ours w/ class distinction +1K 74.06$\pm$2.64 96.23$\pm$0.98 96.74$\pm$1.93 95.72$\pm$0.60
Ours w/o var description +1K 76.06$\pm$4.56 96.29$\pm$1.40 96.41$\pm$2.76 96.18$\pm$1.06
Ours +1K 88.71$\pm$1.98 92.93$\pm$0.91 86.41$\pm$1.85 99.44$\pm$0.27
Figure 4: The difference between Cramér’s V correlation matrices of real and synthetic datasets for categorical variables in the Travel dataset. More intense colors indicate larger differences, with positive differences shown in red and negative differences shown in blue. Black indicates where the correlation is not measured since only one unique value is generated for those variables.
3.3 Exploring optimal prompt design through ablation studies
We investigate key prompt design elements focusing on ML performance and generation efficiency, as detailed in Tables 3 and Table 4. Token efficiency is measured by the number of input tokens required for a fixed number of input samples. LLM generation efficiency is evaluated based on the average number of samples generated per iteration and the success rate of generating at least one valid data sample per inference. This evaluation is based on 100 iterations, each with 20 random input samples.
Our results indicate that CSV-style prompting generally outperforms sentence-style prompting in ML performance with the same number of input samples. While generating data for one class at a time yields good results for the Sick and Income datasets, it results in a low F1 score of 58.67% on the Travel dataset. However, generating data for multiple classes simultaneously with balanced samples improves this score to 70.37%. Furthermore, applying unique variable mapping consistently enhances F1 scores on the Sick dataset. The combined approach of random mapping and grouping achieves higher F1 score of 91.95% compared to 86.60% with random mapping but without grouping, underscoring the positive impact of this combined approach. These findings indicate that prompt design significantly influences the quality of generated data, emphasizing the need for carefully tailored prompts to generate high-quality synthetic tabular data.
For the Travel and Income datasets, performance without grouping is slightly better than with grouping, but the differences are within 0.5%p, making their performance on par. However, these show significant differences in LLM generation efficiency. Without grouping, an average of 0.88 samples are generated per iteration, whereas grouping increases this to an average of 8.6 samples per iteration. Repeating the set twice within the prompt further improves generation efficiency, boosting the average output to over 10 samples and the success rate to 94%, by creating input patterns that LLMs can mimic to generate data. Unique variable mapping also significantly enhances LLM generation efficiency and success rates, as detailed in Appendix B.
Furthermore, we conduct an ablation study on the task specification elements, as shown in Table 5. Our approach of triggering completion significantly enhances overall ML performance compared to the instruction-based prompting methods. Interestingly, adding class distinctions within the prompt achieves the best sensitivity, indicating that the distinction facilitates better generation for minority classes. Providing variable descriptions improves balanced accuracy but reduces the F1 score, suggesting that this element may be optional rather than a definitive improvement. For a comparison of generation efficiency using these methods, see Appendix B.
3.4 Analysis on feature correlation
We compare the feature correlations of the categorical variables between original and synthetic data, as illustrated in Fig. 4. Our method, utilizing Mistral and Llama2, produces data that most closely mirrors the original feature correlations for both minor and major classes. Moreover, when compared to the ablated versions, our grouping method significantly contributes to preserving feature correlation for the minority class. For additional feature correlation results on the Sick dataset, see Appendix C.
4 Conclusion
This study demonstrates the effectiveness of using LLMs for tabular data generation and introduces an optimal solution for prompt design to generate realistic and reliable data without additional training, particularly addressing class imbalance. Our extensive evaluations demonstrate the robustness and effectiveness of our method, which significantly improves ML classification performance on the six real-world tabular datasets. This highlights its practicality and potential impact on improving model performance in various application scenarios, contributing significantly to the field.
Appendix
This supplementary material enhances the main manuscript by providing detailed experimental results and additional visualizations.
  • Appendix A presents more experimental results on the machine learning classification for imbalanced classes.
  • Appendix B offers complete additional ablation study results.
  • Appendix C provides feature correlation results for the Sick dataset.
  • Appendix D studies the impact of sample size on the machine learning performance.
  • Appendix E comprehensively analyzes the experiments conducted on the toy dataset.
  • Appendix F discusses about the broader impacts of our research.
  • Appendix G outlines limitations of our work.
  • Section H provides related work.
  • Appendix I provides comprehensive information on the datasets, comparison baselines, implementation details, and prompt design and examples.
  • Appendix J presents confusion matrix results, highlighting the effectiveness of our method in addressing minor class problems.
  • Appendix K shows the complete results of Table 1.
Appendix A Additional experimental results
Our experiments with synthetic tabular data for imbalanced classes follow three main approaches: (1) Adding synthetic data to both minor and major classes of the existing dataset and evaluating machine learning classification performance. The results of this approach are presented in Table 1 of the main manuscript. (2) Adding synthetic data only to the minor class, with detailed results provided in Appendix A.1. (3) Evaluating classification performance using only synthetic data without the original dataset, as discussed in Appendix A.2.
Table 6: Comparison of classification performance for imbalanced data by generating synthetic data for the minority class to match the size of the majority class. The average performances of the Gradient Boosting Classifier over five runs are reported.
Dataset Method #syn F1 score $\uparrow$ BAL ACC $\uparrow$ Sensitivity $\uparrow$ Specificity $\uparrow$
Travel Original - 60.00$\pm$0.00 72.31$\pm$0.00 60.00$\pm$0.00 84.62$\pm$0.00
+SMOTE +163 63.45$\pm$1.05 75.08$\pm$0.69 68.00$\pm$0.00 82.15$\pm$1.38
+SMOTENC +163 62.61$\pm$2.54 74.74$\pm$2.00 70.40$\pm$3.58 79.08$\pm$0.84
+TVAE +163 54.55$\pm$0.00 68.46$\pm$0.00 60.00$\pm$0.00 76.92$\pm$0.00
+CopulaGAN +163 52.46$\pm$0.00 66.62$\pm$0.00 64.00$\pm$0.00 69.23$\pm$0.00
+CTGAN +163 57.63$\pm$0.00 70.92$\pm$0.00 68.00$\pm$0.00 73.85$\pm$0.00
+GReaT +163 62.07$\pm$0.00 74.46$\pm$0.00 72.00$\pm$0.00 76.92$\pm$0.00
+Ours +163 72.13$\pm$0.00 83.23$\pm$0.00 88.00$\pm$0.00 78.46$\pm$0.00
Income Original - 69.05$\pm$0.00 78.02$\pm$0.00 61.03$\pm$0.00 95.00$\pm$0.00
+SMOTE +13,487 69.62$\pm$0.25 79.84$\pm$0.23 68.86$\pm$0.57 90.81$\pm$0.14
+SMOTENC +13,487 69.79$\pm$0.11 80.51$\pm$0.06 71.80$\pm$0.23 89.22$\pm$0.19
+TVAE +13,487 68.26$\pm$0.00 79.12$\pm$0.00 68.43$\pm$0.00 89.82$\pm$0.00
+CopulaGAN +13,487 67.76$\pm$0.02 81.12$\pm$0.01 80.15$\pm$0.03 82.09$\pm$0.00
+CTGAN +13,487 68.51$\pm$0.00 78.82$\pm$0.00 66.26$\pm$0.00 91.38$\pm$0.00
+GReaT +13,487 70.04$\pm$0.00 83.36$\pm$0.00 85.08$\pm$0.00 81.64$\pm$0.00
+Ours +13,487 71.09$\pm$0.02 83.60$\pm$0.01 83.53$\pm$0.03 83.66$\pm$0.00
A.1 Comparative analysis with adding the new dataset only to the minor class
One of the major advantages of our method is its ability to generate high-quality synthetic data for minor classes, even with limited samples. Similar to oversampling techniques like SMOTE, we add synthetic dataset only to the minor class and evaluate machine learning classification performance. To balance the ratio of major to minor classes, we add only minor class samples to the original data, then measure the classification performance of the machine learning model. For this task, we additionally investigate the effectiveness of our method compared to SMOTE and SMOTENC [6]. As shown in Table 6, adding minor class samples generated from our method leads to higher F1 scores and balanced accuracy on both Travel and Income datasets, compared to when samples generated by other baselines are added. In the Travel dataset, where the number of minor class samples is substantially low, baseline models, especially, CopulaGAN and CTGAN, demonstrate markedly low sensitivity, indicating their failure to accurately learn the minor class data distribution. In contrast, our method demonstrates its efficacy in producing high-quality data for minor classes on the Travel dataset. Compared to the Original, it significantly improves the F1 score by 12.13%p, underscoring its superior performance of our approach.
A.2 Comparative analysis of classification performance using only synthetic data
Table 7: Comparison of classification performance using only synthetic data for training. We report average values for five runs employing Gradient Boosting Classifier.
Dataset Method #syn F1 score $\uparrow$ BAL ACC $\uparrow$ Sensitivity $\uparrow$ Specificity $\uparrow$
Travel Original - 60.00$\pm$0.00 72.31$\pm$0.00 60.00$\pm$0.00 84.62$\pm$0.00
TVAE 1K 46.51$\pm$0.00 63.85$\pm$0.00 40.00$\pm$0.00 87.69$\pm$0.00
CopulaGAN 1K 0.00$\pm$0.00 49.23$\pm$0.00 0.00$\pm$0.00 98.46$\pm$0.00
CTGAN 1K 12.50$\pm$0.00 50.15$\pm$0.00 8.00$\pm$0.00 92.31$\pm$0.00
GReaT 1K 62.50$\pm$0.00 73.85$\pm$0.00 60.00$\pm$0.00 87.69$\pm$0.00
Ours 1K 70.00$\pm$0.00 81.23$\pm$0.00 84.00$\pm$0.00 78.46$\pm$0.00
HELOC Original - 71.32$\pm$0.04 73.39$\pm$0.03 68.47$\pm$0.06 78.31$\pm$0.00
TVAE 1K 68.91$\pm$0.10 71.50$\pm$0.08 65.28$\pm$0.12 77.72$\pm$0.07
CopulaGAN 1K 69.50$\pm$0.00 70.82$\pm$0.00 69.17$\pm$0.00 72.47$\pm$0.00
CTGAN 1K 68.71$\pm$0.02 71.98$\pm$0.03 63.36$\pm$0.00 80.60$\pm$0.05
GReaT 1K 64.14$\pm$0.10 69.54$\pm$0.07 55.76$\pm$0.15 83.33$\pm$0.11
Ours 1K 71.48$\pm$0.06 70.17$\pm$0.05 78.90$\pm$0.16 61.44$\pm$0.16
We compare the classification performance using only synthetic data for training, as shown in Table 7. For both the Travel and HELOC datasets, the classification model trained on synthetic datasets generated by our model outperform the one trained on original data in terms of F1 score and Sensitivity. This improvement in performance suggests that the data generated by our method accurately represents the distribution of the original data. Furthermore, it highlights the effectiveness of our class-balanced generation approach in improving the performance of classification models.
Table 8: Comparison of token efficiency and generation stability for ablating prompt design elements on the Sick dataset
Format Class Balance Group Unique #set Input tokens $\downarrow$ Output samples $\uparrow$ Success rate $\uparrow$
Sentence single - 3867.2 0.52 52%
CSV-style single - 1223.6 2.42 38%
Sentence multi - 3879.9 0.52 52%
CSV-style multi - 1226.3 3.05 48%
Sentence multi - 3870.0 0.63 63%
CSV-style multi - 1117.3 0.70 16%
CSV-style multi - 1985.4 3.00 15%
CSV-style multi 1 1331.3 4.73 64%
CSV-style multi 2 1439.9 6.93 99%
CSV-style multi 2 2060.8 17.68 95%
Table 9: Comparison of token efficiency and generation stability for task specification in prompt design on the Sick dataset.
Method Input tokens $\downarrow$ Output samples $\uparrow$ Success rate $\uparrow$
Instruction-$\text{CuratedLLM}^{*}$ 4677.4 21.67 95%
Instruction-$\text{LITO}^{*}$ 1277.4 4.06 57%
Ours w/ class distinction $2$.7 9.7 91%
Ours w/o var description 2066.9 15.96 88%
Ours 2060.8 17.68 95%
Appendix B Additional ablation study results
This section presents additional ablation study results on the Sick datasets. We evaluate the token efficiency and generation stability of ablated versions based on 100 iterations, each with 20 random input samples. As shown in Table 8, for the Sick dataset, due to its repetitive monotonous values, the number of generated samples is low, even with grouping. Using a unique variable mapper in such cases significantly increases the number of samples generated per attempt from 6.93 to 17.68. This improvement demonstrates that our approach effectively controls the quantity of generated data samples, ensuring the stable production of synthetic tabular data. Moreover, the ability to generate multiple data samples in a single inference significantly can lowers generation time and costs.
We also compare the efficiency and stability of generation when ablating task specification elements in our prompt design, as shown in Table 9. The prompt design from CuratedLLM [25], despite requiring the highest number of input tokens, achieves high performance in terms of output samples and success rate. However, the generated data shows relatively low quality, resulting in a low classification performance with an F1 score and balanced accuracy, as discussed in Section 3.3. In contrast, the prompt design of LITO [31] is the most efficient in terms of input tokens but produces the fewest output samples and has the lowest success rate, indicating instability in synthetic data generation.
Our proposed method balances efficiency and performance, using approximately half the input tokens compared to CuratedLLM’s prompt while generating a high number of output samples and maintaining a high success rate. Additionally, the synthetic data generated by our method achieves the highest classification performance in terms of F1 score, as shown in Section 3.3. When class distinctions are provided, the average number of output samples decreased from 17.68 to 9.70. This result indicates that it is more robust to distinguish groups with simple numbering, such as A and B, rather than explicitly specifying class information in each group. Including variable descriptions slightly reduces the number of output samples and success rate compared to not including them, but the difference is not significant. This suggests that the grouping strategy plays a more crucial role in robust data generation than the presence of variable descriptions.
Figure 5: The differences between Pearson correlation matrices of real data and synthetic data for numerical variables in the Sick dataset. More intense colors indicate larger differences, with positive differences shown in red and negative differences shown in blue.
Appendix C Analysis of feature correlation between numerical variables
In the main manuscript, we visualize the feature correlation between categorical variables. Here, we analyze the Pearson correlation between numerical variables in the Sick dataset and compare the differences in feature correlation values to the original data, as shown in Fig. 5. As a result, our methodology consistently shows feature correlations similar to the original data for both minor and major classes. In contrast, other baselines either show significant deviations in feature correlation for all classes or perform well for major classes but not for minor ones. These results demonstrate that our methodology shows superior performance compared to the baselines in generating high-quality data that accurately represents both minor and major classes within the original dataset.
Appendix D Impacts of sample size on the classification performance
Figure 6: Impact of sample size on the classification performance on the Income and Diabetes datasets when classifiers are trained exclusively on synthetic samples generated from our proposed method. Black dashed lines denote the F1 score using only the original data.
Figure 7: Complete generation results on an imbalanced toy dataset. Major class is shown in blue and minor class in red. Our approach via in-context learning of LLM achieves (1) clear class boundaries, (2) accurate feature correlations, (3) proper value ranges, (4) well-represented relationships between categorical (last row) and numerical variables, and (5) better diversity compared to the fine-tuned GReaT [4] model.
We investigate how varying the number of synthetic samples affects classification performance. This analysis is conducted on the Income and Diabetes datasets, and the results are shown in Fig. 6. For the Income dataset, sample sizes range from 500 to 10,000, while for Diabetes, the range was 500 to 20,000. For both datasets, a consistent increase in the F1 score is observed as the number of samples grow. This result demonstrates that our method can adequately represent the original data distribution by generating a sufficient number of synthetic data samples.
Figure 8: The illustration of the input samples (blue and orange) and the generated samples (green and red) for each inference on the imbalanced toy dataset when the our proposed approach received 180 input samples. Empty spaces in the visualization indicate instances where the model failed to generate samples. To produce a total of 1000 samples, the model performed 23 inferences.
Figure 9: The illustration of the input samples (blue and orange) and the generated samples (green and red) for each inference on the imbalanced toy dataset when the ablated model without grouping received 180 input samples. Empty spaces in the visualization indicate instances where the model failed to generate samples. To produce a total of 1,000 samples, the model performed 37 inferences.
Appendix E Complete qualitative results on toy dataset
In Fig. 7, we illustrate the complete qualitative results on the toy dataset. We assume a challenging multivariate data scenario in an imbalanced binary classification situation, where variables with and without correlations are mixed. Using the scikit-learn library  [22], we created a total of 12 variables. These variables have strong correlations in pairs. There are 11 numerical variables and 1 categorical variable, and the categorical variable is shown on the x-axis in the last row of Fig. 7. In the figure, the major class is shown in blue, and the minor class is shown in red.
The generated results can be evaluated from four aspects. Firstly, whether each data sample is correctly generated to belong to each class, i.e., whether the class boundary is clear. Secondly, whether the shapes formed by the correlation between the variables, such as circles or moons, are well represented. Thirdly, whether the values generated fall within the acceptable range for each variable. Finally, whether the relationship between the categorical and numerical variables is well represented.
The baseline method, GReaT [4], has successfully learned from the training data to generate class-separated samples. However, upon closer inspection, there are samples belonging to different classes. Ours, despite generating samples without training, shows similar or better quality.
Overall, CSV-style generation produces better class separation boundaries compared to sentence-style generation. Regarding class presentation, giving a single class can sometimes result in good generation outcomes, but it fails to accurately capture the correlation with the major class, resulting in shapes like squares instead of circles, and it also fails to accurately generate the categorical variable. On the other hand, when generating multi-class samples, if the sample numbers are not balanced, the generation results for the minor class can be compromised. Even when balanced, providing grouping leads to clearer class distinctions. Additionally, adding unique variable mapping significantly improves the performance, resulting in a well-distinguished categorical variable.
In addition, we visualize the inputs and outputs of both our method and the ablated model without grouping, as shown in Fig. 8 and Fig. 9. For each inference, 180 class-balanced samples are input into the LLM prompt. To generate a total of 1,000 samples, our method required 23 inferences, whereas the ablated model needed 37 inferences, demonstrating the stability of our method in generating synthetic data samples. In terms of data generation quality, our method successfully generates both major and minor classes in most inferences, with the distribution of the generated samples closely mirroring that of the actual input data. In contrast, the ablated model exhibits a higher number of inferences where the generated data deviates from the input data distribution.
Appendix F Broader Impact
Our method is generally applicable to any tabular classification task with minimal pre-processing, making it accessible for various research and industrial applications. The potential positive impacts include enhanced data accessibility and privacy preservation, which can improve decision-making processes in fields like healthcare and finance. Additionally, it reduces the need for using real data, thereby protecting individual privacy. However, there are potential negative impacts, such as the risk of data misuse. Synthetic data generation could be misused for malicious purposes, such as creating fake datasets to deceive systems or individuals.
Figure 10: The illustration of the input (blue and orange) and generated (green and red) samples from our method where input samples are exclusively taken from either the left or right side of the circle.
Appendix G Limitations
If the input samples do not adequately represent the data, the LLM may fail to generate samples that follow the original data distribution, leading to degraded performance. As shown in Fig. 10, if the input samples do not fully capture the original data shape, the generated data may fail to accurately follow this shape. A limitation of our work is that our current method relies on providing a large number of data samples to ensure sufficient representation in in-context learning. Future work will focus on increasing efficiency by selectively choosing key samples that can help the LLM generate data effectively with fewer examples.
Appendix H Related work
In the field of tabular data synthesis, traditional methods have predominantly focused on sampling techniques, such as SMOTE [6], which have been foundational in addressing class imbalances in tabular data. Although useful, these methods often fall short in capturing complex inter-feature correlations [2] and typically lack the versatility needed for regression tasks or addressing multiple feature imbalances. Recently, deep-learning-based generative models have demonstrated their efficacy in generating realistic tabular data [7, 9, 15, 30, 33, 34]. For example, CTAB-GAN [33] introduced a generator that mitigates the problem caused by imbalanced discrete variables. Additionally, diffusion models have also exhibited promising results in tabular data generation. TabDDPM [18] introduced a diffusion process for both numerical and categorical features in tabular data for accurate data generation. Large language model (LLM)-based methods have emerged as a recent trend in tabular dataset generation. GReaT [4] surpassed previous methods by fine-tuning LLMs to generate tabular data formatted in natural text. However, this fine-tuning process is resource-intensive, requiring extensive training of large models for each dataset.
As an alternative, leveraging the in-context learning capabilities of LLMs presents a promising solution. LLMs have demonstrated the ability to solve various problems, learn and mimic patterns from given input data, and generate new data without additional training for parameter updates [5, 13, 17, 19, 28, 29, 32]. Despite this potential, there is limited research on what LLMs generate for synthetic tabular data and what the optimal prompts for high-quality tabular data generation would be. Previous studies [31, 25] have shown the potential of LLMs for in-context learning of tabular data generation. However, these studies have not sufficiently explored optimal prompts for synthetic tabular data generation, leading to instability in the generated samples. This instability necessitates additional validation modules and a two-stage structure.
Given that the performance of in-context learning with LLMs heavily depends on prompt design [10, 26, 28, 32], carefully designing prompts specifically for imbalanced tabular data is essential to maximize LLM capabilities in effectively generating realistic data. Recognizing the importance of optimal prompt design and the potential instability in generated data, we propose a novel method for generating high-quality synthetic tabular data through tailored prompting. The proposed method is generally applicable to various types of tabular datasets without the need for training, thereby offering a cost-effective solution in data synthesis.
Appendix I Additional experimental details
This section provides the illustration of ablated versions (Appendix I.1), comprehensive information on the datasets (Appendix I.2) and comparison baselines (Appendix I.3), implementation details (Appendix I.4), and final prompt design and examples (Appendix I.5).
I.1 Ablation prompt examples
This section visualizes five ablated versions of our proposed prompt design, as shown in Fig. 11. The ablations involve removing or altering four key components: data format, balanced class ratio, multi-class generation, and class grouping.
Figure 11: Illustration of ablated versions.
I.2 Dataset
The detailed description of the datasets are shown in Table 10. The variable descriptions are directly sourced from dataset platforms like Kaggle or the UCI repository. In our approach, we designate the class with fewer instances as the positive class for binary classification tasks.
Table 10: Details of the datasets used to evaluate synthetic tabular data generation models.
Dataset #Class #categorical features #numerical features #samples Domain
Travel 2 4 2 447 Marketing
Sick 2 21 6 3,711 Medical
HELOC 2 - 23 9,872 Financial
Income 2 8 6 32,561 Social
Diabetes 3 36 11 101,766 Medical
Thyroid 2 16 1 364 Medical
I.3 Baseline reproducibility
In our research, we utilize implementations from the Synthetic Data Vault [21] for TVAE [30], CTGAN [30], and CopulaGAN, while for the GReaT [4] model, we employ code from its official repository. In our experiment, the GReaT model fails to accurately generating column names ‘max_glu_serum’ and ‘A1Cresult’ in the Diabetes dataset, even after training for 85 epochs that reported in the original paper. Extending the training to 105 epochs does not rectify this issue, with ‘max_glu_serum’ and ‘A1Cresult’ being correctly generated in only approximately 2.8% and 12.1% of instances, respectively. For these columns, failure cases are treated as NaN value.
We utilize the official open-source code from the official TabDDPM GitHub repository to reproduce CTAB-GAN [33], CTAB-GAN+ [34], and TabDDPM [18]. For the Adult Income and California datasets, the hyperparameters provided are used. For other datasets lacking provided hyperparameters, we apply the hyperparameters from the Wilt dataset, which, according to the corresponding paper, demonstrated the largest performance improvement over the original data.
Due to the lack of official code provided by the authors of LITO [31], we adopt the prompt design described in their original paper. Specifically, we use the prompt for LITO-C to generate minor class data. To generate samples for both majority and minority classes, we also use the modified prompt to generate majority class data. Similarly, for CuratedLLM [25], we follow the prompt design detailed in their paper. Although LITO and CuratedLLM propose multi-stage techniques for data generation, we focus solely on evaluating their prompt designs.
I.4 Experimental setup
For classification, we utilize the target class of each dataset to divide groups. Specifically, the Diabetes dataset has three groups, and the Travel, Sick, HELOC, Income and Thyroid datasets have two groups. For the process of unique variable mapping, each discrete value is transformed into a combination of three characters, including uppercase letters and digits.
Regarding hyperparameter settings of downstream machine learning models: For the Gradient Boosting Classifier, we use the default hyperparameters provided by scikit-learn library. For XGBoost [8], CatBoost [23], and LightGBM [16], we conduct 5-fold cross-validation on the training set of the original data, optimizing two key hyperparameters: learning rate and max depth. These optimized settings are then consistently applied across all experiments, aligning the model’s performance closely with the characteristics of the original data. Where possible, we train the classification models using a single NVIDIA GeForce RTX 3090 GPU on an Ubuntu 18.04 system.
I.5 More details on the prompt design
This section provides an in-depth explanation of the construction of the proposed prompt. Prompt examples for generating synthetic data are provided in Tables 11-16.
Altogether, the final prompt template for synthetic data generation consists of four parts.
Descriptions (Optional). The prompt may start with descriptions of the variables in the dataset, provided line-by-line, if precise and available. These descriptions define what each column in the dataset represents, offering contextual understanding to the LLM about the data. It will be located at the top of the prompt (but not repeated to maintain token efficiency) for additional context.
Set-level examples. Then, a set-level examples are present. This section begins with a data entry format header, specifying the order and names of each feature. This header acts as a guide for the LLM, indicating the start of an example set, which is crucial for interpreting the structure of the tabular data. Following the header, there’s a line specifying a group name, signaling the start of a specific group. Instances belonging to this group are then listed, one per line, in a CSV style. This process is repeated for each group within the set.
Repetition of set-Level template. To reinforce the data structure and patterns, this set-level template is repeated several times within the prompt, each time with distinct data samples. This repetition is key in enabling the LLM to recognize, learn, and subsequently replicate the set format in its synthetic data generation process.
Trigger for LLM data generation. After all the example sets are presented, the same header that indicates the start of an example set is included at the end of the prompt. This header serves as a signal for the LLM to commence generating synthetic data, maintaining the structure and patterns established in the prompt.
By following this structured approach, the LLM is guided to generate synthetic data that not only adheres to the format and structure of the original dataset but also incorporates variability and distinctiveness in the data patterns. The inclusion of descriptions, careful structuring of set-level examples, and strategic repetition of templates all contribute to the effectiveness of this synthetic data generation method.
Appendix J Confusion matrix results
This section provides confusion matrix results on four binary classification datasets with Gradient Boosting classifier. The confusion matrix in Fig. 12 reveals that machine learning classification model classifies most accurately when the data generated from our method is added to the original data.
Figure 12: Confusion matrix on the Travel, Sick, HELOC, and Income datasets using Gradient Boosting classifier.
Table 11: Example of a data synthesis prompt for the Travel dataset. This prompt illustrates the structure of sets and groups. Each iteration utilizes randomly selected samples from the training data.
Template Prompt sample Descriptions Churn: whether customer churns or doesnt churn for tour and travels company,
age: the age of customer,
FrequentFlyer: whether customer takes frequent flights,
AnnualIncomeClass: class of annual income of user,
ServicesOpted: number of times services opted during recent years,
AccountSyncedToSocialMedia: whether company account of user synchronised to their social media,
[…]
Set Header Churn,Age,FrequentFlyer,AnnualIncomeClass,ServicesOpted, AccountSyncedToSocialMedia,BookedHotelOrNot Group A.
Churn,28,Yes,High Income,6,No,Yes
Churn,37,Yes,Low Income,4,Yes,Yes
[…]
Group B.
Doesnt churn,38,No,Low Income,1,Yes,No
Doesnt churn,28,No Record,Low Income,5,No,Yes
[…]
Set Header Churn,Age,FrequentFlyer,AnnualIncomeClass,ServicesOpted, AccountSyncedToSocialMedia,BookedHotelOrNot Group A.
Churn,30,Yes,High Income,4,No,No
Churn,28,No,Low Income,6,No,Yes
[…]
Group B.
Doesnt churn,37,Yes,Low Income,1,No,No
Doesnt churn,36,No,Middle Income,1,No,Yes
[…]
Set Header Churn,Age,FrequentFlyer,AnnualIncomeClass,ServicesOpted, AccountSyncedToSocialMedia,BookedHotelOrNot Group A.
Churn,27,Yes,High Income,5,No,No
Churn,37,No,Low Income,5,Yes,No
[…]
Group B.
Doesnt churn,37,Yes,Low Income,1,No,No
Doesnt churn,36,No,Middle Income,1,No,Yes
[…]
Set Header (trigger) Churn,Age,FrequentFlyer,AnnualIncomeClass,ServicesOpted, AccountSyncedToSocialMedia,BookedHotelOrNot
Table 12: Example of a data synthesis prompt for the Sick dataset. This prompt illustrates the structure of sets and groups. Each iteration utilizes randomly selected samples from the training data. In this scenario, our unique variable mapping is employed, whereby each unique value of a variable is consistently substituted with a unique three-character alphanumeric string. This approach ensures diversity and robustness in the synthesized data.
Template Prompt sample Descriptions Class: hypothyroidism is a condition in which the thyroid gland is underperforming or producing too little thyroid hormone,
age: the age of an patient,
sex: the biological sex of an patient,
TSH: thyroid stimulating hormone,
T3: triiodothyronine hormone,
TT4: total levothyroxine hormone,
T4U: levothyroxine hormone uptake,
FTI: free levothyroxine hormone index,
[…]
Set Header Class,age,sex,on_thyroxine,query_on_thyroxine, on_antithyroid_medication,sick,pregnant,thyroid_surgery, I131_treatment,query_hypothyroid,query_hyperthyroid,lithium, goitre,tumor,hypopituitary,psych,TSH_measured,TSH,T3_measured,T3, TT4_measured,TT4,T4U_measured,T4U,FTI_measured,FTI,referral_source Group A.
[…]
Group B.
[…]
Set Header Class,age,sex,on_thyroxine,query_on_thyroxine, on_antithyroid_medication,sick,pregnant,thyroid_surgery, I131_treatment,query_hypothyroid,query_hyperthyroid,lithium, goitre,tumor,hypopituitary,psych,TSH_measured,TSH,T3_measured,T3, TT4_measured,TT4,T4U_measured,T4U,FTI_measured,FTI,referral_source Group A.
[…]
Group B.
[…]
Set Header (trigger) Class,age,sex,on_thyroxine,query_on_thyroxine, on_antithyroid_medication,sick,pregnant,thyroid_surgery, I131_treatment,query_hypothyroid,query_hyperthyroid,lithium, goitre,tumor,hypopituitary,psych,TSH_measured,TSH,T3_measured,T3, TT4_measured,TT4,T4U_measured,T4U,FTI_measured,FTI,referral_source
Table 13: Example of a data synthesis prompt for the HELOC dataset. This prompt illustrates the structure of sets and groups. Each iteration utilizes randomly selected samples from the training data.
Template Prompt sample Descriptions Set Header RiskPerformance,ExternalRiskEstimate,MSinceOldestTradeOpen, MSinceMostRecentTradeOpen,AverageMInFile,NumSatisfactoryTrades, NumTrades60Ever2DerogPubRec,NumTrades90Ever2DerogPubRec, PercentTradesNeverDelq,MSinceMostRecentDelq, MaxDelq2PublicRecLast12M,MaxDelqEver,NumTotalTrades, NumTradesOpeninLast12M,PercentInstallTrades, MSinceMostRecentInqexcl7days,NumInqLast6M,NumInqLast6Mexcl7days, NetFractionRevolvingBurden,NetFractionInstallBurden, NumRevolvingTradesWBalance,NumInstallTradesWBalance, NumBank2NatlTradesWHighUtilization,PercentTradesWBalance Group A.
Bad,90,211,6,102,17,0,0,100,-7,7,8,17,1,0,0,1,1,0,-8,0,-8,0,0
[…]
Group B.
[…]
Set Header RiskPerformance,ExternalRiskEstimate,MSinceOldestTradeOpen, MSinceMostRecentTradeOpen,AverageMInFile,NumSatisfactoryTrades, NumTrades60Ever2DerogPubRec,NumTrades90Ever2DerogPubRec, PercentTradesNeverDelq,MSinceMostRecentDelq, MaxDelq2PublicRecLast12M,MaxDelqEver,NumTotalTrades, NumTradesOpeninLast12M,PercentInstallTrades, MSinceMostRecentInqexcl7days,NumInqLast6M,NumInqLast6Mexcl7days, NetFractionRevolvingBurden,NetFractionInstallBurden, NumRevolvingTradesWBalance,NumInstallTradesWBalance, NumBank2NatlTradesWHighUtilization,PercentTradesWBalance Group A.
Bad,63,150,3,50,29,3,3,91,75,6,3,32,4,75,-7,6,6,33,90,1,7,1,73
[…]
Group B.
[…]
Set Header (trigger) RiskPerformance,ExternalRiskEstimate,MSinceOldestTradeOpen, MSinceMostRecentTradeOpen,AverageMInFile,NumSatisfactoryTrades, NumTrades60Ever2DerogPubRec,NumTrades90Ever2DerogPubRec, PercentTradesNeverDelq,MSinceMostRecentDelq, MaxDelq2PublicRecLast12M,MaxDelqEver,NumTotalTrades, NumTradesOpeninLast12M,PercentInstallTrades, MSinceMostRecentInqexcl7days,NumInqLast6M,NumInqLast6Mexcl7days, NetFractionRevolvingBurden,NetFractionInstallBurden, NumRevolvingTradesWBalance,NumInstallTradesWBalance, NumBank2NatlTradesWHighUtilization,PercentTradesWBalance
Table 14: Example of a data synthesis prompt for the Income dataset. This prompt illustrates the structure of sets and groups. Each iteration utilizes randomly selected samples from the training data.
Template Prompt sample Descriptions Set Header income,age,workclass,fnlwgt,education,education_num, marital-status,occupation,relationship,race,sex,capital-gain, capital-loss,hours-per-week,native-country Group A.
>50K,52,Private,298215,Bachelors,13,Married-civ-spouse, Craft-repair,Husband,White,Male,0,0,50,United-States
>50K,30,Self-emp-inc,321990,Masters,14,Married-civ-spouse, Exec-managerial,Husband,White,Male,15024,0,60,?
[…]
Group B.
<=50K,42,Private,572751,Preschool,1,Married-civ-spouse, Craft-repair,Husband,White,Male,0,0,40,Nicaragua
<=50K,25,Self-emp-not-inc,159909,Assoc-voc,11,Married-civ-spouse, Farming-fishing,Husband,White,Male,0,0,40,United-States
[…]
Set Header income,age,workclass,fnlwgt,education,education_num, marital-status,occupation,relationship,race,sex,capital-gain, capital-loss,hours-per-week,native-country Group A.
>50K,38,Private,58108,Bachelors,13,Married-civ-spouse, Exec-managerial,Husband,White,Male,0,0,50,United-States
>50K,31,State-gov,124020,Assoc-acdm,12,Married-civ-spouse, Tech-support,Husband,White,Male,0,0,40,United-States
[…]
Group B.
<=50K,51,Private,138514,Assoc-voc,11,Divorced,Tech-support, Unmarried,Black,Female,0,0,48,United-States
<=50K,18,Private,205218,11th,7,Never-married,Sales,Own-child,White, Female,0,0,20,United-States
[…]
Set Header (trigger) income,age,workclass,fnlwgt,education,education_num, marital-status,occupation,relationship,race,sex,capital-gain, capital-loss,hours-per-week,native-country
Table 15: Example of a data synthesis prompt for the Diabetes dataset. This prompt illustrates the structure of sets and groups. Each iteration utilizes randomly selected samples from the training data. In this scenario, our unique variable mapping is employed, whereby each unique value of a variable is consistently substituted with a unique three-character alphanumeric string. This approach ensures diversity and robustness in the synthesized data.
Template Prompt sample Descriptions Set Header readmitted,encounter_id,patient_nbr,race,gender,age,weight, admission_type_id,discharge_disposition_id,admission_source_id, time_in_hospital,payer_code,medical_specialty,num_lab_procedures, num_procedures,num_medications,number_outpatient,number_emergency, number_inpatient,diag_1,diag_2,diag_3,number_diagnoses, max_glu_serum,A1Cresult,metformin,repaglinide,nateglinide, chlorpropamide,glimepiride,acetohexamide,glipizide,glyburide, tolbutamide,pioglitazone,rosiglitazone,acarbose,miglitol, troglitazone,tolazamide,insulin,glyburide-metformin, glipizide-metformin,glimepiride-pioglitazone, metformin-rosiglitazone,metformin-pioglitazone,change,diabetesMed Group A.
AD6,39850434,428733,HX2,LL7,10F,H7H,2,7,1,13,QTD,3S2,44,4,15,1,0, 6,QDY,HR7,8K9,9,JAT,N0H,K06,2TV,65H,A7C,MK7,JWO,HXK,SVQ,BMY,FCV, ZRU,IDJ,8CO,78A,NXL,CWD,6TQ,VDM,HMH,0OU,HNM,NGY,NC4
[…]
Group B.
YRB,163732050,91571517,7MT,LL7,BM0,H7H,3,1,1,1,QTD,SKI,17,6,9,0,0, 1,6RA,4UR,0TR,9,JAT,N0H,K06,2TV,65H,A7C,MK7,JWO,HXK,SVQ,BMY,FCV, ZRU,IDJ,8CO,78A,NXL,PDO,6TQ,VDM,HMH,0OU,HNM,UXM,R51
[…]
Group C.
PS6,274193592,68737500,7MT,LL7,10F,H7H,1,3,7,5,QBR,SKI,60,1,18,0,0, 0,DL8,AL8,P2T,9,JAT,ZAX,K06,2TV,65H,A7C,MK7,JWO,HXK,SVQ,BMY,FCV, ZRU,IDJ,8CO,78A,NXL,CWD,6TQ,VDM,HMH,0OU,HNM,NGY,NC4
[…]
Table 16: Example of a data synthesis prompt for the Thyroid dataset. This prompt illustrates the structure of sets and groups. Each iteration utilizes randomly selected samples from the training data.
Template Prompt sample Descriptions Set Header Recurred,Age,Gender,Smoking,Hx Smoking,Hx Radiothreapy, Thyroid Function,Physical Examination,Adenopathy,Pathology, Focality,Risk,T,N,M,Stage,Response Group A.
Yes,46,M,Yes,No,No,Euthyroid,Single nodular goiter-left, Bilateral,Follicular,Uni-Focal,High,T4b,N1b,M1,II,Structural Incomplete
Yes,27,M,No,No,No,Euthyroid,Multinodular goiter,Bilateral, Papillary,Multi-Focal,Intermediate,T3a,N1b,M0,I, Structural Incomplete
[…]
Group B.
[…]
Set Header Recurred,Age,Gender,Smoking,Hx Smoking,Hx Radiothreapy, Thyroid Function,Physical Examination,Adenopathy,Pathology, Focality,Risk,T,N,M,Stage,Response Group A.
Yes,37,M,No,No,No,Euthyroid,Multinodular goiter,Bilateral, Papillary,Multi-Focal,Intermediate,T3a,N1b,M0,I,Structural Incomplete
Yes,63,M,Yes,No,No,Euthyroid,Single nodular goiter-right, Right,Papillary,Multi-Focal,Intermediate,T3a,N1b,M0,II,Structural Incomplete
[…]
Group B.
[…]
Set Header (trigger) Recurred,Age,Gender,Smoking,Hx Smoking,Hx Radiothreapy, Thyroid Function,Physical Examination,Adenopathy,Pathology, Focality,Risk,T,N,M,Stage,Response
Appendix K Complete results on comparison with baseline methods
Table K offers the complete performance comparison results with standard deviation values of 20 independent experiments on all six datasets.
Table 17: Complete results on comparison of classification performance improvement using synthetic data from baseline models and our method, averaged across four classifiers: XGBoost, CatBoost, LightGBM, and Gradient Boosting classifier, with each model run five times. “#syn” represents the number of synthetic samples added to the original data.
Dataset Method #syn F1 score $\uparrow$ BAL ACC $\uparrow$ Sensitivity $\uparrow$ Specificity $\uparrow$
Travel Original - 58.12$\pm$2.04 (0.00) 71.00$\pm$1.41 (0.00) 57.00$\pm$3.40 (0.00) 85.00$\pm$0.68 (0.00)
+TVAE [30] +1K 59.78$\pm$4.89 (+1.66) 72.35$\pm$3.52 (+1.35) 62.00$\pm$6.81 (+5.00) 82.69$\pm$1.31 (-2.31)
+CopulaGAN [21] +1K 21.76$\pm$2.00 (-36.36) 55.52$\pm$0.83 (-15.48) 12.80$\pm$1.64 (-44.20) 98.23$\pm$1.89 (+13.23)
+CTGAN [30] +1K 29.84$\pm$3.61 (-28.28) 57.79$\pm$1.56 (-13.21) 19.20$\pm$2.46 (-37.80) 96.38$\pm$1.15 (+11.38)
+CTAB-GAN [33] +1K 56.07$\pm$8.29 (-2.05) 69.58$\pm$5.14 (-1.42) 51.00$\pm$7.88 (-6.00) 88.15$\pm$3.31 (+3.15)
+CTAB-GAN+ [34] +1K 54.66$\pm$3.64 (-3.46) 68.62$\pm$2.40 (-2.38) 53.00$\pm$3.40 (-4.00) 84.23$\pm$2.05 (-0.77)
+GReaT [4] +1K 60.95$\pm$2.59 (+2.83) 72.86$\pm$1.80 (+1.86) 58.80$\pm$3.69 (+1.80) 86.92$\pm$0.79 (+1.92)
+TabDDPM [18] +1K 53.20$\pm$4.10 (-4.92) 67.70$\pm$2.69 (-3.30) 50.40$\pm$4.19 (-6.60) 85.00$\pm$1.31 (0.00)
 +Ours +1K  66.65$\pm$2.53 (+8.53) 78.23$\pm$2.10 (+7.23) 78.00$\pm$4.59 (+21.00) 78.46$\pm$2.50 (-6.54)
Sick Original - 87.81$\pm$2.46 (0.00) 91.22$\pm$0.95 (0.00) 82.83$\pm$1.71 (0.00) 99.61$\pm$0.22 (0.00)
+TVAE [30] +1K 87.77$\pm$2.88 (-0.04) 91.47$\pm$1.33 (+0.25) 83.37$\pm$2.47 (+0.54) 99.56$\pm$0.23 (-0.05)
+CopulaGAN [21] +1K 83.60$\pm$1.33 (-4.21) 86.61$\pm$0.53 (-4.61) 73.37$\pm$0.97 (-9.46) 99.86$\pm$0.10 (+0.25)
+CTGAN [30] +1K 87.52$\pm$1.68 (-0.29) 89.86$\pm$0.52 (-1.36) 79.89$\pm$0.97 (-2.94) 99.82$\pm$0.19 (+0.21)
+CTAB-GAN [33] +1K 86.12$\pm$1.49 (-1.69) 89.51$\pm$1.24 (-1.71) 79.35$\pm$2.49 (-3.48) 99.68$\pm$0.06 (+0.07)
+CTAB-GAN+ [34] +1K 82.35$\pm$4.16 (-5.46) 86.28$\pm$2.49 (-4.94) 72.83$\pm$4.86 (-10.00) 99.74$\pm$0.14 (+0.13)
+GReaT [4] +1K 87.23$\pm$1.87 (-0.58) 90.83$\pm$1.09 (-0.39) 82.07$\pm$2.10 (-0.76) 99.60$\pm$0.12 (-0.01)
+TabDDPM [18] +1K 85.17$\pm$2.07 (-2.64) 89.30$\pm$1.18 (-1.92) 79.02$\pm$2.26 (-3.81) 99.57$\pm$0.10 (-0.04)
 +Ours +1K  88.71$\pm$1.98 (+0.90) 92.93$\pm$0.91 (+1.71) 86.41$\pm$1.85 (+3.58) 99.44$\pm$0.27 (-0.17)
HELOC Original - 71.01$\pm$0.47 (0.00) 73.21$\pm$0.31 (0.00) 67.89$\pm$0.82 (0.00) 78.52$\pm$0.34 (0.00)
+TVAE [30] +1K 71.12$\pm$0.32 (+0.11) 73.25$\pm$0.33 (+0.04) 68.15$\pm$0.21 (+0.26) 78.34$\pm$0.48 (-0.18)
+CopulaGAN [21] +1K 71.23$\pm$0.27 (+0.22) 73.32$\pm$0.25 (+0.11) 68.37$\pm$0.31 (+0.48) 78.26$\pm$0.36 (-0.26)
+CTGAN [30] +1K 70.82$\pm$0.24 (-0.19) 73.06$\pm$0.27 (-0.15) 67.60$\pm$0.25 (-0.29) 78.52$\pm$0.55 (0.00)
+CTAB-GAN [33] +1K 70.60$\pm$0.41 (-0.41) 72.87$\pm$0.21 (-0.34) 67.39$\pm$0.97 (-0.50) 78.36$\pm$0.74 (-0.16)
+CTAB-GAN+ [34] +1K 71.03$\pm$0.05 (+0.02) 73.15$\pm$0.10 (-0.06) 68.13$\pm$0.21 (+0.24) 78.17$\pm$0.38 (-0.35)
+GReaT [4] +1K 70.35$\pm$0.33 (-0.66) 72.96$\pm$0.24 (-0.25) 66.22$\pm$0.49 (-1.67) 79.70$\pm$0.21 (+1.18)
+TabDDPM [18] +1K 70.65$\pm$0.18 (-0.36) 72.89$\pm$0.14 (-0.32) 67.51$\pm$0.35 (-0.38) 78.26$\pm$0.30 (-0.26)
 +Ours +1K  71.92$\pm$0.11 (+0.91) 73.66$\pm$0.17 (+0.45) 69.96$\pm$0.21 (+2.07) 77.35$\pm$0.51 (-1.17)
Income Original - 66.90$\pm$2.12 (0.00) 76.45$\pm$1.48 (0.00) 57.28$\pm$3.41 (0.00) 95.61$\pm$0.46 (0.00)
+TVAE [30] +20K 66.96$\pm$1.36 (+0.06) 76.80$\pm$1.11 (+0.35) 59.13$\pm$2.92 (+1.85) 94.48$\pm$0.71 (-1.13)
+CopulaGAN [21] +20K 66.75$\pm$1.72 (-0.15) 76.73$\pm$1.33 (+0.28) 59.16$\pm$3.41 (+1.88) 94.29$\pm$0.81 (-1.32)
+CTGAN [30] +20K 66.22$\pm$0.82 (-0.68) 76.27$\pm$0.65 (-0.18) 57.95$\pm$1.73 (+0.67) 94.59$\pm$0.44 (-1.02)
+CTAB-GAN [33] +20K 66.48$\pm$2.11 (-0.42) 76.31$\pm$1.51 (-0.14) 57.45$\pm$3.61 (+0.17) 95.17$\pm$0.58 (-0.44)
+CTAB-GAN+ [34] +20K 66.49$\pm$1.14 (-0.41) 76.42$\pm$0.90 (-0.03) 58.14$\pm$2.41 (+0.86) 94.70$\pm$0.63 (-0.91)
+GReaT [4] +20K 67.95$\pm$1.36 (+1.05) 77.51$\pm$1.05 (+1.06) 60.69$\pm$2.56 (+3.41) 94.33$\pm$0.51 (-1.28)
+TabDDPM [18] +20K 66.85$\pm$1.83 (-0.05) 76.50$\pm$1.34 (+0.05) 57.70$\pm$3.26 (+0.42) 95.30$\pm$0.60 (-0.31)
 +Ours +20K  69.16$\pm$1.01 (+2.26) 79.15$\pm$0.82 (+2.70) 66.45$\pm$1.98 (+9.17) 91.85$\pm$0.49 (-3.76)
Diabetes Original - 54.87$\pm$1.37 (0.00) 42.07$\pm$1.23 (0.00) 60.00$\pm$0.64 (0.00) 60.73$\pm$1.63 (0.00)
+TVAE [30] +10K 54.79$\pm$1.40 (-0.08) 41.96$\pm$1.24 (-0.11) 59.96$\pm$0.66 (-0.04) 60.71$\pm$1.65 (-0.02)
+CopulaGAN [21] +10K 54.27$\pm$1.48 (-0.60) 41.59$\pm$1.26 (-0.48) 59.73$\pm$0.75 (-0.27) 59.97$\pm$1.65 (-0.76)
+CTGAN [30] +10K 54.72$\pm$1.13 (-0.15) 41.92$\pm$1.02 (-0.15) 59.86$\pm$0.53 (-0.14) 60.63$\pm$1.32 (-0.10)
+CTAB-GAN [33] +10K 54.22$\pm$1.28 (-0.65) 41.53$\pm$1.09 (-0.54) 59.73$\pm$0.59 (-0.27) 59.91$\pm$1.46 (-0.82)
+CTAB-GAN+ [34] +10K 54.24$\pm$1.16 (-0.63) 41.52$\pm$1.01 (-0.55) 59.63$\pm$0.57 (-0.37) 60.01$\pm$1.31 (-0.72)
+GReaT [4] +10K 54.78$\pm$1.31 (-0.09) 41.98$\pm$1.17 (-0.09) 59.98$\pm$0.60 (-0.02) 60.61$\pm$1.55 (-0.12)
+TabDDPM [18] +10K 54.64$\pm$1.50 (-0.23) 41.83$\pm$1.28 (-0.24) 59.91$\pm$0.75 (-0.09) 60.55$\pm$1.78 (-0.18)
 +Ours +10K  54.94$\pm$1.43 (+0.07) 42.14$\pm$1.29 (+0.07) 60.04$\pm$0.66 (+0.04) 60.82$\pm$1.68 (+0.09)
Thyroid Original - 94.23$\pm$1.99 (0.00) 95.08$\pm$1.60 (0.00) 91.14$\pm$3.12 (0.00) 99.02$\pm$1.01 (0.00)
+TVAE [30] +1K 90.45$\pm$1.89 (-3.78) 92.20$\pm$1.65 (-2.88) 86.36$\pm$3.30 (-4.78) 98.04$\pm$0.00 (-0.98)
+CopulaGAN [21] +1K 86.73$\pm$3.99 (-7.50) 88.71$\pm$2.91 (-6.37) 78.41$\pm$5.08 (-12.73) 99.02$\pm$1.01 (0.00)
+CTGAN [30] +1K 76.40$\pm$6.58 (-17.83) 81.14$\pm$4.55 (-13.94) 62.27$\pm$9.10 (-28.87) 100.0$\pm$0.00 (+0.98)
+CTAB-GAN [33] +1K 53.49$\pm$3.71 (-40.74) 68.30$\pm$1.73 (-26.78) 36.59$\pm$3.45 (-54.55) 100.0$\pm$0.00 (+0.98)
+CTAB-GAN+ [34] +1K 27.46$\pm$7.80 (-66.77) 58.07$\pm$2.60 (-37.01) 16.14$\pm$5.21 (-75.00) 100.0$\pm$0.00 (+0.98)
+GReaT [4] +1K 91.31$\pm$1.61 (-2.92) 92.46$\pm$0.99 (-2.62) 85.91$\pm$1.40 (-5.23) 99.02$\pm$1.01 (0.00)
+TabDDPM [18] +1K 94.39$\pm$1.09 (+0.16) 96.26$\pm$0.50 (+1.18) 95.45$\pm$0.00 (+4.31) 97.06$\pm$1.01 (-1.96)
 +Ours +1K  94.80$\pm$1.02 (+0.57) 96.39$\pm$0.62 (+1.31) 95.23$\pm$1.02 (+4.09) 97.55$\pm$0.87 (-1.47)